One side of a triangle is 33 cm shorter than the base. The other side is 73 cm longer than the base. What length of the base will give a perimeter of 223 cm?
Answer by ikleyn(52790)

.
One side of a triangle is 33 cm shorter than the base. The other side is 73 cm longer than the base.
What length of the base will give a perimeter of 223 cm?
~~~~~~~~~~~~~~~~~~~

Let x be the length of the base (now unknown).


According to the condition,


    x + (x-33) + (x+73) = 223    (the perimeter).


Simplify and solve for x:


    3x = 223 + 33 - 73


    3x = 183  ====>  x = 183/3 = 61.


====>  The base should be 61 cm long.


Check.  You should find other side lengths and check if the thriangle inequalities are satisfied.


        The other sides are 61-33 = 28  and  61 + 73 = 134.


        61 + 28 = 89  is less than  134  ==================>  Such a triangle DOES NOT EXIST.

Answer.   SUCH  A  TRIANGLE  DOES  NOT  EXIST.